OpenStudy (anonymous):

What is the value of x in the equation below? 1/2(x + 4) + 3 = 2x( 9/2+3/4 )

OpenStudy (anonymous):

(1/2)x+2+3=2x(21/4) (1/2)x+5=(21/2)x subtrat (1/2)x on both sides 5=10x divide both sides by 10 x=(1/2)

OpenStudy (anonymous):

The answer choices are A. 5/7 B. 1 C. 2 D. 1/2

OpenStudy (anonymous):

Thank you guys ;)